### v2

# MQuant: Unleashing the Inference Potential of Multimodal Large Language Models via Static Quantization

Abstract.
Multimodal large language models (MLLMs) have garnered widespread attention due to their ability to understand multimodal input. However, their large parameter sizes and substantial computational demands severely hinder their practical deployment and application. While quantization is an effective way to reduce model size and inference latency, its application to MLLMs remains underexplored. In this paper, we propose MQuant, a post-training quantization (PTQ) framework designed to tackle the unique challenges of multimodal large language models (MLLMs). Conventional quantization often struggles with MLLMs because of (a) high inference latency from large visual token counts, (b) distributional disparities between visual and textual tokens, and (c) extreme outliers introduced by Hadamard-based transformations. To address these issues, MQuant introduces: • Modality-Specific Static Quantization (MSQ), assigning distinct static scales for visual vs. textual tokens; • Attention-Invariant Flexible Switching (AIFS), reordering tokens to preserve casual attention while eliminating expensive token-wise scale computations; • Rotation Magnitude Suppression (RMS), mitigating weight outliers arising from online Hadamard rotations. On five mainstream MLLMs (including Qwen-VL, MiniCPM-V, CogVLM2), MQuant under W4A8 achieves near-floating-point accuracy (<1% degradation) while reducing inference latency by up to 30%, significantly outperforming existing PTQ baselines. Our MQuant effectively bridges the gap for efficient and accurate MLLMs inference in resource-constrained devices. Code is released in https://github.com/StiphyJay/MQuant.
Multimodal Large Language Model, Model Compression, Model Efficiency, Post-Training Quantization
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755433isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Natural language generationccs: Computing methodologies Visual content-based indexing and retrieval
1. Introduction
Figure 1. (a) Prefill visual tokens counts across different MLLMs as the image splits/resolution increases. (b) The activation values of visual tokens range from 20-20 to $10$, while textual tokens concentrate near 0 (rarely exceeding $\pm 0.5$.
Recent advances in large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023a, b; Dubey et al., 2024) have led to remarkable performance on a wide range of natural language processing tasks. However, these models often struggle when dealing with non-textual data such as images or videos. Multimodal large language models (MLLMs) (Reid et al., 2024; Achiam et al., 2023a; Wang et al., 2023) address this limitation by integrating visual and textual modalities, thereby enabling comprehensive understanding and reasoning. Despite these benefits, their large parameter sizes, coupled with substantial computational demands, poses a major challenge for real-world deployment, particularly in resource-constrained or privacy-sensitive environments.
Challenges in MLLMs Quantization. Quantization has proven an effective strategy for reducing memory usage and inference costs in LLMs (Yuan et al., 2023a; Shao et al., 2023; Zhou et al., 2025), by converting high-precision parameters (e.g., FP32) into lower-bit representations (e.g., INT8). Yet, the transition from LLMs to MLLMs brings three unique difficulties: Time-to-First-Token (TTFT) Explosion. MLLMs often generate large numbers of visual tokens (e.g., patch embeddings or region proposals) with the resolution and aspect ratio of input images or videos. As shown in Fig 1(a), in models such as Qwen2-VL (Wang et al., 2024a), the number of prefill visual tokens grows as image resolution increases (detailed figure in Appendix A.6). As image or video resolution increases, the prefill visual tokens can escalate dramatically—leading to high TTFT and negatively impacting latency-sensitive tasks. Moreover, in higher-demand scenarios such as video-based tasks and multi-image dialogues, the accumulation of visual tokens becomes even more pronounced, further exacerbating TTFT growth. Although the commonly adopted per-token dynamic quantization (Bai et al., 2023b; Wang et al., 2024a; Liu et al., 2024; Chen et al., 2024c) offers flexibility to accommodate the larger activation distribution variance across different tokens in vision modal, its requirement for online token-wise scale recalculation during inference significantly amplifies this computational overhead. Moreover, recent studies (Chen et al., 2024a; Tan et al., 2024a) have underscored the inefficiencies of per-token dynamic quantization, especially for resource-constrained edge devices, where such methods exacerbate latency and memory pressures. Disparate Modality Distributions. As shown in Fig 1 (b), the activation distributions between visual and textual tokens reveal substantial numerical discrepancies, where visual token activations can span a significantly broader range (e.g., 20-20 to $10$) than textual tokens, which typically center near 0. A single global scale factor for both visual and textual tokens leads to either aggressive clipping of visual outliers or increased quantization granularity for text, harming overall accuracy. Visual Token Outliers. High-magnitude outliers in visual tokens, often reflecting salient image regions, make traditional clipping-based methods unsuitable. As shown in Table 1, naive clipping outliers significantly degrades accuracy, highlighting the necessity for judicious outlier treatment.
Table 1. Ablations on the clip range for different tokens.
Clip Range Bits (W/A) Visual Tokens Textual Tokens
- BF16 / BF16 61.40
(0-1.0) BF16 / INT16 61.20 ($\downarrow$0.20) 61.25 ($\downarrow$0.15)
(0-0.99999) BF16 / INT16 18.92 ($\downarrow$42.48) 60.09 ($\downarrow$1.31)
To this end, we propose MQuant, an accurate and efficient post-training quantization (PTQ) solution specifically designed for multimodal large language models (MLLMs). First, to reduce the TTFT while maintaining accuracy, we introduce Modality-Specific Static Quantization (MSQ), which apply distinct static per-tensor scales for visual vs. textual tokens based on the distinct distribution differences. Second, for further acceleration, we design an Attention-Invariant Flexible Switching (AIFS) scheme. AIFS transforms mixed multimodal tokens into unified, modality-decoupled tokens without the need for dynamic position vectors, maintaining computational equivalence and memory efficiency by avoiding the overhead associated with dynamic processing. Based on MSQ and AIFS, MQuant slashes TTFT and preserves accuracy. Third, we reveal the weight outliers caused by the online Hadamard rotations through theoretical analysis and propose Rotation Magnitude Suppression (RMS) to mitigate them. Our theoretical analysis reveals the emergence of large mean channels in MLLM weights, and RMS effectively reduces these outliers with minimal overhead.
We evaluate MQuant on five mainstream MLLMs, including InternVL (Chen et al., 2024b), Qwen-VL (Bai et al., 2023b), MiniCPM-V (Yao et al., 2024), CogVLM2 (Hong et al., 2024) and Qwen2-VL (Wang et al., 2024a). The extensive results on diverse multimodal reasoning tasks demonstrate that MQuant achieves less than 1% accuracy loss on all MLLMs with 23% and 100% speedup in prefill and decode stage under the W4A8 setting, highlighting its superior performance and wide practical value. Our main contributions are summarized as follows:
  • Insightful Observation. To our best knowledge, we present the first comprehensive analysis of quantization issues in MLLMs, revealing the root causes of performance collapse and identifying inference speed bottlenecks as well as the quantization challenges posed by modality differences.
  • MSQ and AIFS. We design Modality-specific Static Quantization (MSQ) and Attention-Invariant Flexible Switching (AIFS) to accelerate inference for heterogeneous variable-sequence multimodal inputs while maintaining accuracy.
  • RMS. We identify weight outliers caused by online rotation, and propose Rotation Magnitude Suppression to enhance quantization performance.
  • MQuant. We propose MQuant, a general PTQ framework designed for MLLMs, demonstrating both near-lossless performance and with significant speedup.
2. Preliminaries and Related Work
2.1. Multimodal Large Language Models
Figure 2. MLLM’s architecture.
As shown in the Fig. 2, the existing MLLMs framework  (Bai et al., 2023b; Chen et al., 2024c) mainly consists of three modules: (1) Visual encoder $\bm{E}$ for processing visual inputs, (2) Vision-language projector $\bm{P}$ for aligning the text and visual modalities, (3) Large language model $LLM$ that handles the multi-modal tokens and performs reasoning.
Vision Encoder.
Taking the input image or video $\bm{X}_{v}$ as input and compressing the original vision information into more compact patch features $\bm{F}_{v}$. This process typically utilizes a Vision Transformer (ViT)  (Dosovitskiy, 2021), such as CLIP (Radford et al., 2021) and OpenCLIP (Ilharco et al., 2021). It can be formulated as: $\bm{F}_{v}=E(\bm{X}_{v})$.
Vision-Language Projector.
The task of the vision-language projector $\bm{P}$ is to map the visual patch features $\bm{F}_{v}$ into the textual feature space: $\bm{E}_{v}=P(\bm{F}_{v})$.
Large Language Model.
The pre-trained large language model is the core component of MLLMs, endowing the framework with exceptional capabilities, such as zero-shot generalization, instruction following, and in-context learning. Typically, a text tokenizer is integrated with the LLM, mapping text prompts $\bm{X}_{t}$ to the text tokens $\bm{E}_{t}$. The text tokens $\bm{E}_{t}$ and the visual tokens $\bm{E}_{v}$ are then concatenated to form the input for MLLMs, which outputs the final response sequence $\bm{O}_{a}$ in an autoregressive manner:
(1) ${LLM}(\bm{O}_{a}|\bm{E}_{v},\bm{E}_{t})=\prod^{l}_{i=1}{LLM}(y_{i}|\bm{E}_{v},\bm{E}_{t},y_{<i})\vskip-2.84526pt$
where $l$ denotes the length of $\bm{O}_{a}$. The parameter sizes of large language models (LLMs) range from 3 billion to tens of billions. Commonly used open-source LLMs include the Llama series (Touvron et al., 2023a, b), Qwen (Bai et al., 2023a), InternLM (Cai et al., 2024), MiniCPM (Hu et al., 2024b), ChatGLM (GLM et al., 2024).
MLLMs’ foundational components have greatly benefited from rapid advancements in LLM technology. Flamingo (Alayrac et al., 2022) pioneered connecting pre-trained visual encoders to LLMs, demonstrating strong generalization across visual-language tasks. Following ChatGPT (Achiam et al., 2023b), numerous open-source models based on pre-trained LLMs, such as LLaMA series (Touvron et al., 2023a, b; Dubey et al., 2024), have been proposed (Li et al., 2023a; Huang et al., 2024; Zhu et al., 2023; Liu et al., 2024). Subsequent efforts, including Qwen-VL (Bai et al., 2023b), InternVL (Chen et al., 2024c), and CogVLMV2 (Hong et al., 2024), have enhanced MLLMs by improving high-resolution inputs and scaling training data. Besides, smaller MLLMs like Mini-Gemini (Li et al., 2024c), MobileVLM (Chu et al., 2024), and MiniCPM-V (Yao et al., 2024) have emerged. Despite those advancements, the large parameter sizes of MLLMs lead to high computational costs, yet dedicated quantization methods to reduce memory usage and accelerate inference still remain underexplored.
2.2. PTQ for LLMs/MLLMs
PTQ (Zhou et al., 2024; Jiang et al., 2024; Yu et al., 2025) serves as a potential strategy for model compression. By converting the high-precision variables of pre-trained models into low-bit integers, it achieves memory reduction and inference speed acceleration. For uniform quantization, given a floating-point (FP) tensor $x$ (weights or activations), it can be uniformly quantized to $b$-bits in signed quantization as follows:
(2) $\hat{\mathbf{x}}=\mathbf{Q}_{U}(\mathbf{x},b)=(clamp(\lfloor\frac{\mathbf{x}}{s}\rceil+z,q_{min},q_{max})-z)\cdot s$
where $s=\frac{\max(|\mathbf{x}|)}{2^{b-1}-1}$ is the scale factor, $\lfloor\cdot\rceil$ is the rounding-to-nearest operator, and the function $clamp(\cdot)$ clips values outside the integer range $\left[q_{min},q_{max}\right]$. $z$ is zero-point. $s$ denotes the quantization scale factor, which reflects the proportional relationship between FP values and integers. $\left[q_{min},q_{max}\right]$ is the quantization range determined by the bit-width $b$. Generally, when we quantize the network’s weight with 4-bit and activations with 8-bit, called it as W4A8. We can calculate $s$ offline using the activations from calibration samples, known as static quantization. We can also use the runtime statistics of activations to get $s$, referred to as dynamic quantization. More details are in Appendix A. 12.
Existing post-training quantization (PTQ) methods for LLMs are categorized into weight-only and weight-activation quantization (Yuan et al., 2024; Li et al., 2023b, 2024b; Yue et al., 2024; Hu et al., 2024a, 2025; Chen et al., [n. d.]; Xu et al., 2025, [n. d.]). Weight-only methods like GPTQ (Frantar et al., 2022), QuIP (Chee et al., 2024), and AWQ (Lin et al., 2023) achieve high compression rates but offer limited inference acceleration. In contrast, weight-activation quantization methods (Xiao et al., 2022; Wei et al., 2022; Yuan et al., 2023b; Zhang et al., 2024) quantize both weights and activations, improving memory usage and latency. The main challenge is activation outliers causing quantization errors. Techniques like SmoothQuant (Xiao et al., 2022) shift quantization difficulty from activations to weights, while OmniQuant (Shao et al., 2023) optimizes performance by training quantization parameters. SliceGPT (Ashkboos et al., 2024a) reduces memory demands by designing a Pre-LN + Rotate Scheme for LLMs sparsification based on computational invariance. Recently, Quarot (Ashkboos et al., 2024b) introduced applying offline Hadamard transforms and a partially online Hadamard transform to eliminate outliers, achieving state-of-the-art quantization results on LLMs. However, this solution is not applicable to MLLMs due to inherent modality differences. Building on the above studies, research on PTQ specifically tailored for MLLMs remains notably scarce. Existing studies, such as Q-VLM (Wang et al., 2024b), propose a block-wise quantization framework that optimizes cross-layer dependencies via entropy-guided partitioning. MBQ (Li et al., 2024a) uses gradients of the supervised fine-tuning loss with respect to vision and language tokens as sensitivity indicators to balance reconstruction loss during calibration. Despite efforts, existing methods still employ per-token dynamic quantization for activation, which inherently introduces a significant computational overhead than per-tensor static quantization discussed earlier. In contrast, our MQuant framework addresses these challenges by extending full static quantization to both weights and activations, specifically designed to unleash the inference potential of MLLMs while overcoming the modality-specific limitations faced by LLM-focused PTQ methods, which are not directly applicable to MLLMs due to significant modality differences.
3. Method
In this section, we present MQuant, a post-training quantization solution specifically designed for MLLMs. In Sec. 3.1, we describe proposed modality-specific static quantization (MSQ) and attention-invariant flexible switching (AIFS) mechanisms. In Sec. 3.2, we identify the weight outliers caused by the online Hadamard rotations and state Rotation Magnitude Suppression (RMS). We provide the detailed MQuant algorithm for FP MLLMs in Appendix A.1 Algorithm 1.
3.1. Modality-Specific Static Quantization and Attention-Invariant Flexible Switching
We address this issue with a two-part method: (1) Modality-Specific Static Quantization (MSQ) and (2) Attention-Invariant Flexible Switching (AIFS). MSQ applies different static scaling factors for visual and textual tokens, while AIFS reorders their positions to avoid irregular data slicing. MSQ+AIFS retains accuracy while eliminating the expensive token-wise scale online computation during inference. This approach achieves efficiency gains by leveraging per-tensor static quantization, especially for high-resolution inputs.
Figure 3. Overview of Modality-Specific Static Quantization (MSQ) and Attention-Invariant Flexible Switching (AIFS). AIFS reorders tokens so that all visual tokens appear first, then the textual tokens, while adjusting the causal mask to preserve the original model logic.
Modality-Specific Static Quantization (MSQ)
Let $E$ be an input sequence of length $L$ that intermixes textual and visual tokens. Denote $E=\{\;e^{t}_{1},\dots,e^{t}_{m-1},\,e^{v}_{m},\dots,e^{v}_{n},\,e^{t}_{n+1},\dots,e^{t}_{L}\}$, where $m,n$ specify the visual token segment. We observe that visual tokens often have larger activation magnitudes, which can overshadow textual features. To handle these differences, we apply two distinct sets of static per-tensor quantization parameters:
(3) $E\;=\;\underbrace{(e^{t}_{1},\dots,e^{t}_{m-1})}_{\textstyle\text{text scale }s_{t}}\;\;\underbrace{(e^{v}_{m},\dots,e^{v}_{n})}_{\textstyle\text{visual scale }s_{v}}\;\;\underbrace{(e^{t}_{n+1},\dots,e^{t}_{L})}_{\textstyle\text{text scale }s_{t}}.$
The scaling factors $s_{v}$ and $s_{t}$ correspond to visual and textual tokens, respectively. Our innovation lies in performing single-shot calibration of these factors before inference, thereby eliminating per-token computational overhead. This per-tensor static quantization strategy provides three distinct advantages: (1) complete avoidance of dynamic scale updates, (2) hardware-friendly implementation benefiting from simplified quantization kernels, and (3) prevention of distribution saturation where high-magnitude visual activations could overwhelm the narrower dynamic range of textual tokens.
Attention-Invariant Flexible Switching (AIFS)
Despite the benefits of MSQ, it complicates data handling if textual and visual tokens remain interleaved in $E$. Naïve slicing and concatenating at runtime can increase memory traffic and reduce efficiency for GEMM-based layers like QK and FC. To overcome this, we propose Attention-Invariant Flexible Switching (AIFS). Figure 3 illustrates the idea: we reorder the sequence so that all visual tokens appear first, followed by textual tokens. We then alter the causal mask to preserve the original auto-regressive relationships. Causal attention (Vaswani et al., 2017) ensures each token can only attend to earlier tokens. Formally, for a naive sequence $E$, the attention matrix is
[…]
With AIFS mechanism, we can reorder $E$ and obtain an unified causal mask $M^{u}_{i,j}$, the new $E^{u}=\{\;e^{v}_{m},\dots,e^{v}_{n},\;e^{t}_{1},\dots,e^{t}_{m-1},\,\dots,e^{t}_{n+1},$ $\dots,e^{t}_{L}\}$, and $M^{u}_{i,j}$ could be formulated as:
[…]
The position embeddings are also shifted accordingly (see Appendix A.2), preserving numerical equivalence of attention scores with the original sequence. This maintains auto-regressive consistency while enabling streamlined memory access: all visual/textual tokens can now be processed through respective static-scale multiplications.
Seamless Integration with FlashAttention
MSQ and AIFS, which optimizes MLLMs through optimizing on input token efficiency, while FlashAttention (Dao et al., 2022) accelerates attention computation via reduced memory access and tiling. These approaches are conceptually orthogonal, suggesting seamless integration with additive benefits without inherent conflicts. Specifically, at the beginning of inference, MSQ+AIFS reorders the intermixed multimodal tokens and passes the original visual modality start/end indices ($m$ and $n$) to FlashAttention to ensure that irrelevant positions are correctly masked. Detailed results are in Table  9.
Efficiency and Practical Benefits
Our experiments in Section 4 show that MSQ + AIFS delivers three main advantages: High Compatibility and Strong Theoretical Equivalence: The causal-mask transformation ensures that the model’s output is numerically the same as if tokens were not reordered. This reordering can be integrated into existing MLLM implementations without major code changes. Reduced Inference Latency: We replace per-token dynamic quantization with a static per-modality approach. This removes the need to recalculate scales per token, cutting runtime overhead and boosting throughput. Enhanced Memory Efficiency: Sequential processing of visual (via $s_{v}$) and textual ($s_{t}$) tokens with dedicated scale factors minimizes memory fragmentation, yielding up to 24.7% speedups and 152.9% memory efficiency gains (Table 17) through eliminated padding/slicing operations.
In summary, MSQ + AIFS offers a straightforward and effective way to handle the unique challenges of multi-modal token distributions. It is valuable on edge devices or other constrained platforms where dynamic quantization is hard to implement at scale, yet per-tensor static quantization is hardware-friendly and fast (Tan et al., 2024b). More discussions is in Appendix A.10
3.2. FHT-induced Weight Outlier Mitigation: Rotation Magnitude Suppression
[…]
(6) $|W_{ij}|\;=\;\bigl{|}\,e_{i}^{\top}W\,e_{j}\bigr{|}\;\;\leq\;\;\mu\,\frac{\|W\|_{F}}{\sqrt{mn}},$
Generally, the smaller $\mu$ is, the easier it is to quantize $W$. They also showed that applying a Hadamard transform to both weights and activations can effectively reduce $\mu$. However, Multi-modal LLMs (MLLMs) combine an LLM component with a visual encoder, which often relies on LayerNorm. Quarot’s partial online transform cannot be applied directly to all such norms. Inspired by SliceGPT (Ashkboos et al., 2024a), we convert visual encoders’ LayerNorms into RMSNorms. More details can be found in the Appendix A.14. This change makes Quarot-like Hadamard transforms applicable to MLLMs. Yet, as Table 3 demonstrates, Quarot still underperforms on many MLLM tasks. Quip# (Tseng et al., 2024) also proved that random Hadamard transforms (RHT) can reduce incoherence for Hessians and weights, but they did not analyze online (fast) Hadamard transforms (FHT). In Quarot, FHT is crucial for low-bit quantization (e.g., 4-bit) as shown in Appendix  A.9. We investigate why FHT can yield fresh outliers in MLLMs and propose a method to suppress them. Besides, due to the difference of MLLMs architecture, we also discuss the rotation-based implementation details before RMS in Appendix  A.13.
Figure 4. (a) The pipeline of Quarot, showing offline and partially online Fast Hadamard transforms. (b) Weight matrix where applying online FHT produces outliers in the first row.
Channel-Mean Outliers in Online FHT
Following (Eq. 6), let $W_{\ell_{2}}\in\mathbb{R}^{n\times m}$ be a weight matrix. A Hadamard matrix $H\in\{-1,1\}^{n\times n}$ with orthonormal rows preserves this norm, i.e., $\|\,H\,W_{\ell_{2}}\|_{F}\\ =\|W_{\ell_{2}}\|_{F}$. Suppose $\mu_{\ell_{2}}$ is the incoherence coefficient for $W_{\ell_{2}}$, satisfying $\max(|W_{\ell_{2}}|)=\mu_{\ell_{2}}\,\frac{\|W_{\ell_{2}}\|_{F}}{\sqrt{mn}}$. Let $HW_{\ell_{2}}$ be the transformed weight, and let its incoherence coefficient be $\mu_{H\ell_{2}}$. Then we have:
(7) $\frac{\mu_{H\ell_{2}}}{\mu_{\ell_{2}}}\;=\;\frac{\max\bigl{|}\,H\,W_{\ell_{2}}\bigr{|}}{\max\bigl{|}\,W_{\ell_{2}}\bigr{|}}\,.$
For many Hadamard transforms, the first row (and first column) contain identical $\tfrac{1}{\sqrt{n}}$ entries, while other rows sum to zero. Hence, the first channel after transformation is
(8) $(HW_{\ell_{2}})_{0j}\;=\;\sqrt{n}\,\mathrm{mean}\bigl{(}w_{:,j}\bigr{)}.$
If the mean is large, the first element in that row can exceed the original maximum and thus raise $\mu_{H\ell_{2}}$. Concretely, when
(9) $\sqrt{n}\,\mathrm{mean}\bigl{(}w_{:,j}\bigr{)}\;>\;\max_{i}\,\bigl{(}w_{ij}\bigr{)},$
a new channel-mean outlier appears in the first row. Figure 4(b) shows such an occurrence in an MLLM. This issue arises especially in Quarot’s online (partial) FHT step, which applies Hadamard rotations per forward pass rather than strictly offline. The detailed approve in Appendix  A.3.
Table 2. Compliance Ratio with Eq. 9 Across Model Components
Model Part All Blocks Meet Eq. 9 Ratio
Number Number
Internvl2-8B Visual 24 24 100%
LLM 32 4 12.5%
Qwenvl-9.6B Visual 48 48 100%
LLM 32 13 40.6%
MiniCPM-V-2.6-8B Visual 27 27 100%
LLM 28 3 10.7%
GLM-4V-9B Visual 63 63 100%
LLM 40 6 15.0%
Qwen2vl-7B Visual 32 32 100%
LLM 28 1 3.5%
Prevalence of Eq. 9 Compliance
We conduct systematic block-level verification across five state-of-the-art MLLMs (InternVL2-8B, QwenVL-9.6B, MiniCPM-V-2.6-8B, GLM-4V-9B, and Qwen2VL-7B) to quantify the occurrence of Fast Hadamard Transform (FHT)-induced weight outliers. As evidenced in table  2, our analysis reveals two critical findings: (1) Universal Presence in Vision Encoders: All visual $down_{proj}$ layers (100%) exhibit FHT-induced weight outliers that satisfy Eq. 9, as conceptually illustrated in Fig. 4(b). (2) Variable Manifestation in LLMs: The phenomenon appears in only 3%–40% of LLM $down_{proj}$ layers across different architectures. This empirical validation confirms above theoretical soundness and the critical need for specialized handling of FHT-generated weight outliers, particularly in vision components where their universal presence fundamentally impacts model quantization.
Rotation Mitigation Scheme (RMS)
To this end, we propose RMS to handle these new outliers with minimal overhead. We first identify whether a channel meets (Eq. 9). If it does, we: Split the problematic channel from the main GEMM kernel and process it using a separate GEMV. Zero out that row in $HW_{\ell_{2}}$ so that the main kernel does not double-count it. Add the partial output (from the split GEMV) back to the main path before the activation. Figure 5 shows the workflow. This targeted split ensures large-mean channels do not trigger extreme first-row values during the per-forward-pass FHT. The added cost is small, since only channels meeting Eq. 9 require this procedure. Detailed RMS algorithm in Appendix A.7 outlines how RMS integrates with Quarot’s FHT. Overall, RMS addresses a key shortcoming of online Hadamard transforms by suppressing large-mean channels.
Figure 5. An overview of our proposed RMS. We separate outlier-prone channels into a dedicated GEMV path, zero their row in the main kernel, and then merge the results.
Why Not Subtract Channel Means Directly?
A straightforward idea is to separate out each channel’s mean before applying $H$ and then re-inject it afterward. However, this leads to two major issues: 1. The separated means are still subject to Hadamard rotation, creating a new linear transformation where the first channel again becomes large. 2. Splitting and then re-injecting the means effectively doubles the linear operations, significantly increasing the computational cost. In contrast, our RMS approach modifies only the row triggering the outlier condition and does not require additional linear layers, thereby offering a more efficient and effective resolution.
4. Experiments
Table 3. Comprehensive quantization results of different MLLMs across various multimodal reasoning datasets.
[…]
Visual LLM
- BF16 BF16 77.65 90.97 794 2209
RTN W8A8 W4A8 52.02 59.04 542 1528
SQ 59.88 59.75 544 1540
Quarot 73.34 84.07 715 2067
InternVL2 MQuant 77.49 90.27 785 2156
-8B RTN W4A8 W4A8 40.06 31.58 302 1482
SQ 46.48 31.21 310 1540
Quarot 49.10 33.62 361 1941
MQuant 76.62 88.42 725 2155
- BF16 BF16 61.40 60.36 493 1834
RTN W8A8 W4A8 0.45 0.03 189 625
SQ 7.45 7.70 160 797
Quarot 45.32 42.44 286 940
Qwen-VL MQuant 61.16 59.31 483 1691
-Chat-9.6B RTN W4A8 W4A8 1.02 0.02 193 585
SQ 8.59 4.28 188 921
Quarot 46.77 37.35 289 1091
MQuant 60.50 58.72 473 1713
- BF16 BF16 79.10 89.18 847 2248
RTN W8A8 W4A8 61.00 65.16 332 1300
SQ 62.40 65.76 424 1510
Quarot 73.71 80.04 736 1850
MiniCPM-V MQuant 80.41 89.15 844 2244
2.6-8B RTN W4A8 W4A8 60.70 62.23 351 1404
SQ 65.67 60.02 455 1491
Quarot 68.96 79.63 685 1734
MQuant 81.14 89.75 839 2189
- BF16 BF16 82.82 81.16 782 2153
RTN W8A8 W4A8 7.05 3.70 0.00 140
SQ 9.05 4.10 0.00 148
Quarot 82.00 80.17 782 2115
GLM-4V MQuant 82.06 80.53 782 2164
-9B RTN W4A8 W4A8 7.61 3.60 0.00 163
SQ 9.85 4.40 0.00 188
Quarot 64.16 45.52 516 2048
MQuant 81.58 79.67 754 2120
- BF16 BF16 84.43 93.87 842 2319
RTN W8A8 W4A8 33.92 52.61 442 1298
SQ 49.11 53.97 444 1500
Quarot 79.36 89.57 754 2045
Qwen2-VL MQuant 84.43 93.61 830 2269
-7B RTN W4A8 W4A8 40.20 38.82 422 1082
SQ 46.25 52.36 411 1535
Quarot 71.44 83.96 670 1911
MQuant 84.32 93.58 824 2255
- BF16 BF16 85.48 95.95 883 2479
RTN W8A8 W4A8 37.21 40.94 426 1444
SQ 50.33 55.41 480 1601
Quarot 80.03 91.21 781 2299
Qwen2-VL MQuant 85.48 95.90 880 2469
-72B RTN W4A8 W4A8 28.21 25.94 426 1137
SQ 47.11 54.95 413 1586
Quarot 71.86 86.11 701 2264
MQuant 85.03 95.49 868 2471
Models and Datasets
We evaluate our MQUANT on five MLLMs: InternVL2-8B (Chen et al., 2024b), Qwen-VL-Chat-9.6B (Bai et al., 2023b), MiniCPM-V 2.6-8B (Yao et al., 2024), Qwen2-VL-7B (Wang et al., 2024a), and GLM-4V-9B (Hong et al., 2024). Evaluations are conducted on four benchmarks covering OCR and general question answering: TextVQA (Singh et al., 2019), DocVQA (Mathew et al., 2021), OCRBench (Liu et al., 2023), and MME (Fu et al., 2023), which assesses perception and cognition across 14 tasks. These MLLMs’ details are in Appendix A.11.
Baselines and Implementation Details
We test W8A8 and W4A8 quantization settings for both visual encoders and LLMs, comparing RTN, SmoothQuant (Xiao et al., 2022), and Quarot (Ashkboos et al., 2024b). Notably, we apply static per-tensor activation quantization for both components, unlike the dynamic per-token quantization typically used in existing MLLMs. The calibration dataset consists of 256 randomly selected samples from the corresponding benchmark training sets (Singh et al., 2019; Mathew et al., 2021; Liu et al., 2023). The batch size for latency measurement is 1.
4.1. Overall Results
As shown in Table 3, MQuant can apply to the quantization of various MLLMs, and demonstrates significant improvements over several representative quantization methods. In W8A8 setting, MQuant achieves near-lossless performance to the FP models across diverse datasets and different model scales (7-72B). Notably, even in the more challenging W4A8 setting, MQuant maintains comparable performance with FP models, while other advanced quantization methods exhibit significant performance degradation. These results indicate that our MQuant provide a general and effective PTQ solution with strong compatibility for maintaining high accuracy in MLLMs under various bits settings across diverse multi-modal reasoning tasks. This suggests MQuant maintains robustness (e.g., stability to typical input or models changes) in these frameworks.
Table 4. Comparison of existing MLLMs Quantization Methods
Method W bits A bits Scale GEMM kernel Extra ops
Q-VLM (Wang et al., 2024b) NF4 INT4 per-channel static BF16 GEMM dequant W&A
QSLAW (Xie et al., 2024) INT4 BF16 weight-only BF16 GEMM dequant W
MBQ (Li et al., 2024a) INT4 INT4 per-token dynamic INT4 GEMM per token scale
MQuant (ours) INT4 INT4 per-modality static INT4 GEMM none
Table 5. ScienceQA (Lu et al., 2022) test Results on LLaVA1.5-13B
[…]
BF16 - 90.19 93.14 87.09 89.39 87.06 89.83 90.00
w4a4 Q-VLM (Wang et al., 2024b) 82.55 73.32 83.18 81.03 70.82 86.74 80.78
w4a4 MQuant 87.07 85.24 87.05 88.49 85.50 89.77 87.18
w4 only QSLAW (Xie et al., 2024) 83.26 91.79 80.00 82.80 81.30 83.07 83.70
w4 only MQuant (g=128) 90.21 91.95 87.05 89.20 86.66 89.94 89.16
Table 6. Performance Comparison under Different Precision and Methods (T.VQA, D.VQA, OCRB, MME) on Qwen2VL-72B.
[…]
BF16 - 85.48 95.95 883 2479
W4A8 MBQ (Li et al., 2024a) 83.41 92.30 854 2402
W4A8 MQuant (ours) 85.03 95.49 868 2471
4.2. Comparison with other MLLMs Quantization Methods
To rigorously evaluate MQuant, we compare it against recent state-of-the-art (SOTA) quantization methods (Wang et al., 2024b; Xie et al., 2024; Li et al., 2024a) specifically designed for MLLMs. Table 4 provides a high-level comparison of the quantization schemes. Existing methods often introduce substantial runtime overhead. For example, MBQ (Li et al., 2024a) employs online dynamic per-token quantization; Q-VLM (Wang et al., 2024b) and QSLAW (Xie et al., 2024) require dequantization and perform GEMM in resource-intensive BF16 precision. In contrast, MQuant is the only method that achieves a fully static, per-modality quantization scheme that is directly compatible with highly efficient INT4 GEMM kernels. This design choice is the foundation of its superior inference speed. Besides, due to their different evaluation benchmark, we compare them on their reported results. The empirical results in Table 5 and Table 6 confirm MQuant’s accuracy advantage. On the ScienceQA benchmark (Lu et al., 2022) (Table 5) in LLaVA1.5-13B (Liu et al., 2024), our W4A4 MQuant surpasses Q-VLM by a significant margin of 6.4% in average accuracy. In the weight-only setting, MQuant outperforms QSLAW by 5.46%, highlighting the effectiveness of our activation quantization strategy. Similarly, on Qwen2VL-72B(Wang et al., 2024a) (Table 6), our W4A8 MQuant achieves near-lossless performance, outperforming MBQ while avoiding its costly per-token dynamic quantization. This superior performance is a direct result of our novel designs. The MSQ effectively handles the distributional disparity between visual and textual tokens, a challenge that other methods do not explicitly address. Furthermore, the AIFS scheme eliminates the need for dynamic scales without compromising model accuracy, enabling the use of fast, hardware-friendly static quantization. Together with RMS, these contributions allow MQuant to achieve both the highest accuracy and the fastest true INT4 inference among existing MLLM quantization methods.
4.3. Ablation Study
In this section, we select Qwen2-VL-7B (Wang et al., 2024a), one of the most powerful open-source MLLMs, to validate the effectiveness of our designs.
Table 7. Ablation study of proposed designs on Qwen2-VL-7B (Wang et al., 2024a) with W4A8 setting.
[…]
Static AIFS + MSQ RMS
BF16 84.43 93.87 842 2319 6523
71.44 83.96 670 1911 5479
78.95 87.55 721 2095 5484
84.32 93.58 824 2255 5471
Ablation Study of Proposed Quantization Designs
We perform ablations on Qwen2-VL-7B (Table 7) to isolate the contributions of each quantization component. Beginning with a baseline that applies GPTQ + Hadamard transformations to both the LLM and vision parts with per-tensor static quantization, we progressively introduce MSQ+AIFS for the LLM. This significantly boosts accuracy and latency, highlighting their effectiveness on multimodal input tokens. Next, incorporating RMS to suppress newly arising outliers achieves near-floating-point performance, underscoring the overall robustness and efficiency of our pipeline. We also demonstrate that the design of RMSis also effective for LLMs in Appendix  A.5.
Table 8. Ablation study about calibration dataset size.
[…]
BF16 84.43 93.87 842 2319
128 84.28 93.50 820 2243
256 84.32 93.58 824 2255
512 84.32 93.57 823 2254
Selection of Calibration Dataset Size
Following  (Xiao et al., 2022; Ashkboos et al., 2024b), we selected a sample subset to get the statistics of activations during the calibration process. This ensures the calibration data reflects the distribution in each dataset. To assess sensitivity to calibration dataset size of MQuant, we conducted experiments on Qwen2-VL-7B, including [128, 256, 512] samples. The Table 8 shows the consistency of MQuant in W4A8 with loessless accuracy. These results, revised in supplementary, demonstrate that MQuant is largely insensitive to calibration dataset size. We believe this robustness enhances MQuant’s practical applicability. Therefore, we select 256 randomly selected samples from the corresponding benchmark training set.
Table 9. Latency combining AIFS with FlashAttention from 100 to 8100 tokens ($Q/K/V$ shape is (1, 28, num-tokens, 128))
Tokens Number 100 900 2500 4900 8100
FlashAttn (ms) 27.3 33.6 35.5 165.9 271.6
AIFS+FlashAttn (ms) 27.3 34.1 36.4 167.5 272.1
Table 10. The accuracy and speedup of our MSQ and AIFS scheme on the linear layer during prefill stage. The input sequence is structured as "text-image-text-image-text-image-text" with an image resolution of $2240\times 2240$ and 50 textual tokens. Latency were tested on an NVIDIA RTX 6000 Ada Generation.
[…]
BF16 BF16 84.43 93.87 842 2319 1.690 -
BF16 W4-g128(AWQ) 83.93 (-0.50) 93.13 (-0.74) 828 (-14) 2252 (-67) 2.057 (+0.367) -17.8%
A8-per-token dyn W4-per-channel sta 84.32 (-0.11) 93.61 (-0.26) 830 (-12) 2269 (-50) 1.253 (-0.437) +34.9%
A8-per-tensor sta 40.20 (-44.12) 38.82 (-54.79) 422 (-408) 1082 (-1187) 1.016 (-0.674) +66.3%
A8-MSQ 84.32 (-0.11) 93.61 (-0.26) 830 (-12) 2269 (-50) 1.085 (-0.605) +55.8%
A8-MSQ+AIFS 84.32 (-0.11) 93.61 (-0.26) 830 (-12) 2269 (-50) 1.017 (-0.673) +66.2%
Table 11. Comparison of latency and memory saving with Pytorch (BF16), AWQ (W4-only) and MQuant (W4A8) on Qwen2-VL-7B. $\downarrow$ means lower values are better, $\uparrow$ means larger values are better.
Image size Latency (s) Memory (G)
[…]
8402 0.261 0.304 (+0.043) -14.14% 0.210 (-0.051) +24.76% 16.45 7.45 (-9.00) +120.67% 6.50 (-9.95) +152.92%
19602 1.369 1.598 (+0.229) -14.26% 1.112 (-0.257) +16.63% 17.82 8.82 (-9.00) +100.00% 7.85 (-9.97) +119.93%
30802 5.208 5.872 (+0.664) -11.27% 4.488 (-0.720) +16.02% 20.58 11.58 (-9.00) +77.60% 10.61 (-9.97) +96.45%
56002 8.380 9.393 (+1.013) -10.78% 7.469 (-0.911) +12.19% 22.22 13.22 (-9.00) +57.54% 12.25 (-9.97) +61.65%
Integrate AIFS with Flash Attention
As evidenced in Table 9, we integrated the causal attention mechanism of AIFS into FlashAttention and tested it on an RTX 6000, observing that its speed is nearly identical to that of standard causal attention. The results demonstrates seamless integration MBQ+AIFS with Flash Attention, exhibiting negligible latency overhead (<0.3% across all tested token lengths). The minimal performance gap (maximum 1.6 ms) stems from two key design choices: (1) AIFS’s token reordering operates exclusively in the preprocessing phase without interfering with core attention computation, and (2) our metadata propagation mechanism (visual token indices m/n) preserves Flash Attention’s native memory access patterns. Notably, the 0.96% latency increase at 4,900 tokens validates strong scalability for high-resolution inputs. This highlight the hardware compliance advantages of MBQ and AIFS designs: CUDA-aligned token layouts enable binary compatibility with FlashAttention kernels, which is particularly effective for real-time processing of high-resolution multimodal inputs.
Figure 6. Illustration the weight distributions for the down-proj layer (block 21) in Qwen2VL-7B’s visual encoder under three conditions: (a) original weight, (b) weights after FHT, and (c) weights after FHT with our RMS. More resutls refer to Appendix A.15
Weight Distribution with Rotation Magnitude Suppression (RMS)
As shown in Fig.  12, we visualizes the magnitude of weights outliers without (w/o) and with (w/) RMS. The weight distribution reveals that FHT amplification increases weight magnitudes, while RMS effectively reduces these magnitudes by a factor of 1.0 to 0.1, significantly improving quantization stability (Other distribution are in Appendix A.15). This is also consistent with our ablation results, shown that RMS significantly enhances performance for both MLLMs (Table  7). Taking Qwen2-VL as example, the vision encoder’s down-projection layer (3584×1280, C_inC_outC\_in*C\_out) and LLM’s corresponding layer (18944×3584) demonstrate the efficiency of our approach. The RMS implementation converts a W4A8 GEMM operation into a combined W4A8 GEMM + W4A8 GEMV computation, where the GEMV is selectively applied only to affected channels. This introduces negligible overhead, precisely 1/3584 and 1/18944 of the total computation for vision and LLMs parts.
Figure 7. The Speedup of AIFS+MSQ on Qwen2-VL-7B.
Accuracy & Speedup of AIFS
While MSQ addresses the modality discrepancy, AIFS further reorders visual and textual tokens into a unified sequence to enable efficient per-tensor static quantization of activations. Table 10 shows that MSQ+AIFS not only achieves the same speedup as naïve per-tensor static quantization but also maintains near-floating-point accuracy across linear layers. Figure 7 further illustrates that AIFS yields speedups of 20%–80% as resolution increases, corroborating that rearranging tokens avoids the high overhead of dynamic per-token quantization.
4.4. Latency and Memory Analysis
We evaluate MQuant’s performance across three key metrics: (1) inference speedup, (2) memory efficiency (tested over image resolutions from $280^{2}$-$5600^{2}$), and (3) decoding acceleration. All experiments employ the established ”text-image-text” input format in (Duan et al., 2024) with a fixed 15 textual tokens unless explicitly varied.
Table 12. Speedup of MSQ + AIFS on W4A8 setting.
Stage BF16 Per-token Dyn. Ours Ours + GEMV Speedup
Prefill 1690 1253 1017 - +23%
Decode 17.5 16.4 13.06 8.2 +100%
Inference Speedup and Memory Savings
Overall Speedup As shown in Table 11, MQuant surpasses BF16 and AWQ (W4-only) across all image resolutions, achieving up to 24.76% speedup over PyTorch at $840\times 840$. Even at higher resolutions (e.g., $5600^{2}$), MQuant maintains a 12.19% latency improvement, demonstrating scalability. Memory Efficiency. Our method provides memory savings over both BF16 and AWQ, with reductions exceeding 100% compared to PyTorch (e.g., 152.92% at $840^{2}$). These benefits primarily arise from (1) eliminating the overhead of token-wise scale computation, and (2) converting mixed tokens into modality-decoupled tokens, avoiding slicing and concatenation when image resolutions become large. More resolution experiments are in Appendix  A.8. Decoding Acceleration. We also measure decode time for generating 2,000 tokens with a custom W4A8 GEMV kernel (Table 12). Compared to per-token dynamic quantization, our AIFS+MSQ framework gains 23% speedup in the prefill stage and 100% speedup in decode stage. By shifting online per-toekn dynamic quantization to an offline per-tensor static approach, we greatly reduce inference overhead, especially for long-sequence tasks. Moreover, since visual tokens are generally pricier than textual ones (Duan et al., 2024), these improvements translate to notable real-world cost reductions (e.g., $\approx 30\%$ for OpenAI token pricing).
Table 13. Multi-Batch speedup comparison of MSQ + AIFS on W4A8 setting. Each row shows the cumulative total of text tokens, images, and textual responses for multi-turn inference.
[…]
Text Image Text bfp16 MQuant bfp16 MQuant bfp16 MQuant
1 10 $\times2240$ 50 2.54 1.93 +31.6% 18.01 12.89 +39.7% 20.55 14.82 +38.7%
2 10/10 $\times2240$ / $\times2240$ 50/100 5.42 4.15 +30.6% 37.82 31.56 +19.8% 43.24 35.71 +21.1%
3 10/10/10 $\times2240$ / $\times2240$ / $\times2240$ 50/100/150 8.24 6.42 +28.3% 48.03 40.35 +19.0% 56.27 46.77 +20.3%
4 10/10/10/10 $\times2240$ / $\times2240$ / $\times2240$ / $\times2240$ 50/100/150/200 11.17 8.67 +28.9% 59.09 49.92 +8.4% 70.26 58.59 +20.0%
Table 14. Latency comparison under multi-turns setting.
Turns Config in one Turn All(s) Improve $\uparrow$
Text Image Text bfp16 Ours
1 10 2240x2240 50 20.55 14.82 +38.7%
2 10 2240x2240 50 44.06 32.61 +35.1%
3 10 2240x2240 50 76.67 59.48 +28.9%
Acceleration for Multi-batch and Multi-turn Inference
For multi-batch scenarios, we use $\times2240$ resolution images with ”text-image-text” inputs, we evaluate batches containing 50-200 text tokens while generating 512 decoded tokens. Sequence alignment employs left-padding (pad_token_id) with proper masking (see Appendix A.4). AIFS maintains full multi-batch compatibility without overhead, enabling consistent 20% faster inference versus FP baselines across batch sizes 1-4 (Table 13). For multi-turn dialogue, under identical resolution and input format (50 text tokens per turn + 512 decoded tokens), MQuant achieves 38.7% faster end-to-end inference than BF16 for 1-3 turns (Table 14) through optimized key-value caches and position IDs across turns.
Table 15. Comparison under weight-only settings on Qwen2-VL-7B (Wang et al., 2024a). $\dagger$} means re-implementation based on the official weight-only setting with a group size of 128.
[…]
Visual LLM
- BF16 BF16 84.43 93.87 842 2319
GPTQ (g128) BF16 W8 84.33 93.97 842 2313
GPTQ (g128) BF16 W4 84.18 93.25 831 2285
AWQ (g128) BF16 W4 83.93 93.13 828 2252
MQuant (g128) BF16 W4 84.55 93.18 832 2304
MQuant (g128) W4 W4 84.70 93.57 828 2292
MQuant W4A8 W4A8 84.32 93.58 824 2255
Weight-only Quantization
We evaluate our weight-only quantization using GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2023) under consistent settings (group size=128), quantizing only the LLM while keeping the visual encoder in BF16. As shown in Table  15, our W4 LLM quantization achieves near-lossless accuracy comparable to existing methods. When extended to W4 quantization, MQuant maintains comparable performance with the BF16 baseline. The W4A8 quantization similarly shows competitive or superior results compared to weight-only approaches. These experiments validate MQuant’s effectiveness across: (1) weight-only vs. weight-activation schemes, and (2) partial vs. full-model quantization scenarios.
5. Conclusion
In this paper, we propose MQuant, an accurate and efficient PTQ framework specifically designed for MLLMs. Our approach addresses the challenges of MLLMs by applying Modality-Specific Static Quantization (MSQ) to handle distribution mismatches between visual and text tokens, and introducing an Attention-Invariant Flexible Switching (AIFS) mechanism to reduce TTFT with static per-tensor scaling. Furthermore, we reveal that online Hadamard rotations can lead to weight outliers, and propose Rotation Magnitude Suppression (RMS) to mitigate this issue. Extensive experiments on 5 mainstream MLLMs demonstrate that MQuant achieves near-floating-point accuracy under the W4A8 setting, highlighting its potential to advance MLLMs quantization and practical deployment in resource-constrained environments. For the quantization of video-based MLLMs, we also discuss in Appendix A.16.
Appendix A Appendix
A.1. MQuant Algorithm
Here, we present overall MQuant algorithm for MLLMs in Algorithm 1.
Algorithm 1 MQuant Quantization Algorithm
Input: Full-precision (FP) MLLM model with a vision encoder $\bm{E}$, visual-language projector $\bm{P}$, and a large language model LLM; Calibration dataset $D^{c}$.
Output:
  • For $\bm{E}$ and $\bm{P}$: per-channel weight scale $s^{E}_{w}$, per-channel weight zero-point $z^{E}_{w}$, per-tensor activation scales $s^{E}_{a}$, per-tensor activation zero-point $z^{E}_{a}$.
  • For LLM: per-channel weight scale $s^{llm}_{w}$, per-channel weight zero-point $z^{llm}_{w}$, per-tensor activation scales $s^{{}^{llm}}_{a^{v}}$ for visual tokens and $s^{llm}_{a^{t}}$ for textual tokens, per-tensor activation zero-points $z^{llm}_{a^{v}}$ for visual tokens and $z^{llm}_{a^{s}}$ for textual tokens.
1: Apply Hadamard Rotation to the LLM Part as described:
2: Apply the offline and online Hadamard rotations to all the weights and activations in LLM.
3: Quantize Weights of the LLM:
4: Input the calibration dataset $D^{c}$ to the FP MLLM.
5: Use GPTQ to quantize the weights for LLM, obtaining per-channel weight quantization parameters $s^{llm}_{w}$ and $z^{llm}_{w}$.
6: For the LLM Part:
  1. (a)
    Input $D^{c}$ to the FP MLLM.
  2. (b)
    Compute per-tensor activation quantization parameters $s^{{}^{llm}}_{a^{v}}$ and $s^{llm}_{a^{t}}$, $z^{llm}_{a^{v}}$ and $z^{llm}_{a^{t}}$ for visual and textual tokens respectively, based on the proposed Modality-Specific Static Quantization (MSQ) in Sec 3.1.
  3. (c)
    Reorder the input mixed token sequence from $E={e^{t}_{1},\dots,e^{v}_{m},\dots,e^{v}_{n},\dots,e^{t}_{L}}$ to a unified modality-decoupled sequence $E^{u}={e^{v}_{m},\dots,e^{v}_{n},e^{t}_{1},\dots,e^{t}_{m-1},\dots,e^{t}_{n+1},\dots,e^{t}_{L}}$ using the proposed Attention-Invariant Flexible Switching (AIFS) scheme in Sec 3.1.
7: Transform all the LayerNorm to RMSNorm in MLLM vision encoder $\bm{E}$ and Visual-Language Projector $\bm{P}$ using the proposed Post LayerNorm-to-RMSNorm transformation.
8: Apply the offline and online Hadamard rotations to all the weights and activations in $\bm{E}$ and $\bm{P}$.
9: Quantize Weights of $\bm{E}$ and $\bm{P}$:
10: Input $D^{c}$ to the transformed FP vision encoder $\bm{E}$ and Visual-Language Projector $\bm{P}$.
11: Use GPTQ to quantize $\bm{E}$ and $\bm{P}$, obtaining per-channel weight quantization parameters $s^{E}_{w}$ and $z^{E}_{w}$.
12: Address the weight outliers using caused by online Hadamard based on the proposed Rotation Magnitude Suppression (RMS) in Sec 3.2.
A.2. Rotary Position Embedding for Attention-Invariant Flexible Switching
Many modern LLMs (Touvron et al., 2023a, b; Dubey et al., 2024) use rotary position embedding (RoPE) (Su et al., 2021) to encode information about the order of tokens in the input sequence. Rotary position embeddings are linear transformations applied to keys and queries defined as:
(10) $R_{\Theta,m}^{d_{h}}=\begin{pmatrix}\cos{i\theta_{1}}&-\sin{i\theta_{1}}&\cdots&0&0\\ \sin{i\theta_{1}}&\cos{i\theta_{1}}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&\cos{i\theta_{d_{h}/2}}&-\sin{i\theta_{d_{h}/2}}\\ 0&0&\cdots&\sin{i\theta_{d_{h}/2}}&\cos{i\theta_{d_{h}/2}}\end{pmatrix}$
where $i\in[1,L]$ is the token index, $\Theta=\{\theta_{i}=10000^{-2(i-1)/D},i\in[1,2,...,D/2]\}$, and $\theta_{i},\,i\in 1..D/2$ are predefined constants.
In the proposed Attention-Invariant Flexible Switching (AIFS) mechanism, we also apply the rearrangement for position embedding to maintain the computation equivalence. For a mixed input token $E=\left\{e^{t}_{1},...,e^{v}_{m},...,e^{v}_{n},...,e^{t}_{L})\in(\bm{E}_{v},\bm{E}_{t})\right\}$ (in Eq. 3), where $m$ and $n$ denote the start and end indices of the visual tokens. Specifically, after AIFS, the unified token can formulated as: $E=\left\{e^{v}_{m},...,e^{v}_{n},e^{t}_{1},...,e^{t}_{m-1},...,e^{t}_{L})\in(\bm{E}_{v},\bm{E}_{t})\right\}$. Therefore, the unified token indices after AIFS can be represented as:
(11) (m,,n,1,,m1,n+1,,L)=AIFS(1,,m,,n,,L)(m,...,n,1,...,{m-1},{n+1},...,L)=AIFS{(1,...,m,...,n,...,L)}
Due to we are aware of the final indices for input token after AIFS, than we can utilize the reorder token indices for Eq 10 to get the corresponding position embedding.
A.3. Weights Outliers after Online Hadamard Transformation
Following (Tseng et al., 2024; Ashkboos et al., 2024b) we make use of fast Hadamard transformation where convenient. Hadamard matrix is an orthogonal matrix with entries proportional to $\{+1,-1\}$. A Walsh-Hadamard matrix is a square matrix of size $2^{n}$ with For a Hadamard matrix:
(14) $\bm{H}_{2}=\tfrac{1}{\sqrt{2}}\left[\begin{array}[]{cc}1&1\\ 1&-1\end{array}\right]\qquad\textrm{and}\qquad\bm{H}_{2^{n}}=\bm{H}_{2}\otimes\bm{H}_{2^{n-1}}\,.$
(15) $(AB)_{ij}=\sum_{r=1}^{n}a_{ir}b_{rj}=a_{i1}b_{1j}+a_{i2}b_{2j}+\cdots+a_{in}b_{nj}\vskip-11.38109pt$
Thereby the $\bm{H}W_{\ell_{2}}$ can be formulated as:
(16) $(\bm{H}W_{\ell_{2}})_{ij}=\sum_{r=1}^{n}h_{ir}w_{rj}=h_{i1}w_{1j}+h_{i2}w_{2j}+\cdots+h_{in}w_{nj}\vskip-11.38109pt$
where $\bm{H}\in\mathbb{R}^{d_{in}\times d_{in}}$ and $W_{\ell_{2}}\in\mathbb{R}^{d_{in}\times d_{out}}$, $d_{in}$ and $d_{out}$ is the dimension of input and output of weight $W_{\ell_{2}}$. Due to the properties of the Hadamard matrix $\bm{H}$, whose first row consists entirely of $1$, for the first row in $(\bm{H}W_{\ell_{2}})$, $(\bm{H}W_{\ell_{2}})_{0j}=\sum_{r=1}^{n}w_{rj}$, due to the property of Hadamard matrix $\bm{H}$. So, the values in $W_{\ell_{2}}$ are subject to continuous accumulation and summation, resulting in the exists of outliers in the first row of the output matrix $\bm{H}W_{\ell_{2}}$. Notably, the matrix $\bm{H}W_{\ell_{2}}\bm{Q}$ still has the same problem, for simplicity, we omit the matrix $\bm{Q}$ in the main paper.
A.4. Attention Mask in AIFS when Multi-batch Inference.
Figure 8. The illustration of causal mask when batch size > 1.
During multi-batch inference, we first identify the longest token length within the batch. Subsequently, we left-pad the shorter sequences with $pad_{t}oken_{i}d$ to align all batches to this maximum length. By applying left padding, the padding tokens are associated with the image modality. Additionally, the padded regions are assigned a mask value of 0, ensuring that they do not interfere with attention computations and thereby do not affect the final results. For clarity, we also plot an illustration of causal mask when batch size >1.
A.5. Effectiveness of Rotational Magnitude Suppression for Weight Outliers in LLMs.
Table 16. WikiText-2 perplexity (PPL) on 4-bit quantized LLaMA2 models with an input sequence length of 2048. Lower PPL is better.
Method Weight Quantization PPL
LLaMA2 7B LLaMA2 13B LLaMA2 70B
Baseline 5.47 4.88 3.32
QuaRot GPTQ 6.10 5.40 3.79
QuaRot + RMS GPTQ 6.04 5.32 3.67
For LLMs, compared to the original Quarot, integrating RMS with Quarot leads to performance improvements across LLaMA2 models with 7B, 13B, and 70B parameters, as detailed in Table. 4. For MLLMs, ablation studies presented in Table 7 demonstrate that the RMS method significantly enhances quantization performance.
Figure 9. The number of prefill visual tokens across different MLLMs as the image splits or resolution increases.
A.6. Image Tokens in Various MLLMs
As shown in Fig 9, for different MLLMs (Bai et al., 2023b; Wang et al., 2024a; Yao et al., 2024; Chen et al., 2024c), the number of prefill visual tokens grows as image resolution increases. This rapid expansion in token count exacerbates the inference latency, particularly in per-token dynamic quantization, which requires separate processing for each token, such as memory access and scale computation. As a result, the TTFT in MLLMs increases drastically, severely impacting overall inference latency. Moreover, in higher-demand scenarios such as video-based tasks and multi-image dialogues, the accumulation of visual tokens becomes even more pronounced, further exacerbating the increase in TTFT.
A.7. RMS Algorithm
Here, we present the algorithm of our RMS design with Quarot in Algorithm 2.
Algorithm 2 RMS Integration with Quarot
0:  An LLM or MLLM model
0:  Quantized model with RMS
1:  Initialize an empty list marksmarks
2:  for each linear layer in the modelmodel do
3:   if Layer satisfies Eq. 9 then
4:    Mark the layer and append its ID to marksmarks
5:   end if
6:  end for
7:  Apply Quarot or other Hadamard-based transformations to the modelmodel
8:  for each layer ID in marksmarks do
9:   Modify the layer’s implementation using the RMS method
10:  end for
11:  Quantize the modelmodel
12:  return modelmodel
Table 17. Comparison of latency and memory saving with Pytorch and AWQ on Qwen2-VL-7B. means the Qwen2-VL-7B official implementation.
Image size Pytorch (BF16) AWQ (W4-only) MQuant (W4A8) H $\times$ W Latency(s) Memory(G) Latency(s) Memory(G) Latency(s) Memory(G) 2802 0.257 16.45 0.286 (-10.14%) 7.45 (+120.67%) 0.220 (+16.82%) 6.50 (+152.92%) 5602 0.252 16.45 0.292 (-13.70%) 7.45 (+120.67%) 0.211 (+19.48%) 6.50 (+152.92%) 8402 0.261 16.45 0.304 (-14.14%) 7.45 (+120.67%) 0.210 (+24.76%) 6.50 (+152.92%) 11202 0.326 16.58 0.384 (-15.10%) 7.56 (+119.51%) 0.262 (+24.48%) 6.61 (+151.59%) 14002 0.559 16.90 0.652 (-14.29%) 7.90 (+113.92%) 0.432 (+20.24%) 6.97 (+142.71%) 16802 0.881 17.33 1.066 (-17.39%) 8.33 (+108.57%) 0.705 (+18.23%) 7.40 (+130.27%) 19602 1.369 17.82 1.598 (-14.26%) 8.82 (+100.00%) 1.112 (+16.63%) 7.85 (+119.93%) 22402 2.013 18.40 2.294 (-12.24%) 9.40 (+95.74%) 1.653 (+17.83%) 8.44 (+117.84%) 25202 2.820 19.04 3.175 (-11.14%) 10.05 (+89.59%) 2.357 (+19.63%) 9.10 (+109.92%) 28802 3.883 19.77 4.345 (-10.64%) 10.77 (+83.81%) 3.297 (+17.69%) 9.82 (+101.07%) 30802 5.208 20.58 5.872 (-11.27%) 11.58 (+77.60%) 4.488 (+16.02%) 10.61 (+96.45%) 33602 6.814 21.46 7.548 (-9.73%) 12.46 (+70.65%) 6.004 (+13.41%) 11.50 (+83.01%) 36402 8.360 22.22 9.377 (-10.84%) 13.22 (+57.54%) 7.469 (+11.91%) 12.25 (+61.65%) 44802 8.349 22.22 9.379 (-10.97%) 13.22 (+57.54%) 7.469 (+11.71%) 12.25 (+61.65%) 56002 8.380 22.22 9.393 (-10.78%) 13.22 (+57.54%) 7.469 (+12.19%) 12.25 (+61.65%)
A.8. Speedup and Memory Savings with Scaling Image Resolution
We fixed the input sequence as "text-image-text" with 15 textual tokens and presented the detailed changes of speed and memory, varying the image resolution from $280\times 280$ to $5600\times 5600$. Notably, the "text-image-text" sequence setting is not arbitrarily chosen; instead, it is a common setting in existing evaluation datasets (Duan et al., 2024). We evaluate speedup and memory savings by comparing PyTorch’s BF16, AWQ (W4-only), and our MQuant (W4A8). Speedup: As shown in Table 17, MQuant consistently achieves speedups over both PyTorch and AWQ across all resolutions, with a maximum of 24.76% over PyTorch at $840\times 840$. Notably, MQuant outperforms AWQ, which is slower than PyTorch at most resolutions due to negative speedups. This significant speedup highlights the advantage of our per-tensor static quantization, eliminating the overhead of token-wise scale computation. Even at higher resolutions (e.g., $5600^{2}$), MQuant maintains a 12.19% latency improvement, demonstrating scalability across various image sizes. Memory Savings: MQuant offers substantial memory reductions compared to both PyTorch and AWQ. It consistently reduces memory usage by over 100% compared to PyTorch (e.g., 152.92% at $840^{2}$) and significantly outperforms AWQ’s memory efficiency, achieving up to 101.07% savings at higher resolutions. These experiments demonstrate MQuant’s strengths in both latency and memory savings, achieving up to 24.76% faster inference and reducing memory consumption by over 100% compared to baseline methods. This makes MQuant a more efficient solution for deploying MLLMs in resource-constrained environments.
A.9. Comparison of Online vs. Without Online Hadamard Transform in Quarot
In this section, we evaluate how online Hadamard transformations affect perplexity (PPL) in Quarot for different LLaMA2 models (7B and 13B) under two bit settings, W4A8KV4 and W4A4KV4. As shown in Table 18, enabling online Hadamard transforms yields consistent improvements, especially under more aggressive quantization (e.g., W4A4KV4).
Table 18. Comparison of perplexity (PPL) with or without online Hadamard transforms in Quarot, evaluated on LLaMA2 models of sizes 7B and 13B. Lower PPL is better.
Model Bit Setting PPL
With Online Without Online
LLaMA2 7B W4A8KV4 5.736 5.830
LLaMA2 7B W4A4KV4 6.358 14.375
LLaMA2 13B W4A8KV4 5.123 5.146
LLaMA2 13B W4A4KV4 5.585 24.401
We observe that online Hadamard transforms provide substantial gains when using W4A4KV4 for both 7B and 13B models, reducing PPL from over 14.3 to 6.36 (7B) and from 24.4 to 5.58 (13B), respectively. These results demonstrate the effectiveness of online Hadamard transformations in maintaining quantization accuracy at lower bit precision.
A.10. Advantage of Proposed MSQ and AIFS
In per-tensor static quantization, the quantization parameters (i.e., scale and zero-point) are precomputed for an entire tensor (e.g., weights or activations) and remain fixed throughout inference. While efficient, this approach often leads to large and unacceptable accuracy loss in MLLMs due to their diverse activation distributions across varying inputs.
In contrast, per-token dynamic quantization computes quantization parameters on-the-fly for each input token during inference. This approach incurs significantly higher computational overhead, as the quantization parameters must be recalculated for every input token, along with multiple additional memory traversals. Such requirements make per-token dynamic quantization unfriendly or impractical for edge devices and some AI accelerators, which struggle with fine-grained dynamic operations (Tan et al., 2024b). This issue is especially severe in MLLMs, where the token count increases significantly with higher image resolution or more video frames. The Modality-Specific Static Quantization (MSQ) in MQuant is a novel per-modality quantization approach specifically designed to address the unique challenges of MLLMs quantization.
Furthermore, MSQ can be naturally applied to the unified modality-decoupled tokens generated by AIFS. By integrating MSQ and AIFS, our designs yields three key advantages: (1) Computational Equivalence and Strong Compatibility: The unified causal mask and token index introduced by AIFS preserves the inherent causal relationships among tokens, ensuring numerical equivalence during attention computations. Moreover, since AIFS requires only a one-time rearrangement of the input data (adjust causal mask and token index in offline), it does not alter the overall computation graph. This characteristic allows for seamless integration with other LLM inference acceleration methods, such as FlashAttention (Dao et al., 2022), ensuring both computational equivalence and strong compatibility. As shown in Table 3, MQuant achieves SOTA quantization performance across 5 mainstream MLLMs. (2) Reduced Inference Latency: MSQ not only addresses the substantial distributional differences between modalities but also mitigates the significant computational overhead and increased inference latency caused by the surge in token counts from higher image and video resolutions. As shown in Table 10, MSQ+AIFS significantly reduces latency from 2.057s to 1.1017s, closely matching the speed of the per-tensor static setting while maintaining near-lossless accuracy comparable to the original Float model. (3) Enhanced Memory and Computational Efficiency: By combining MSQ and AIFS, we convert mixed input tokens into unified, modality-decoupled tokens, eliminating the irregular memory operations (e.g., slice, concat, pad) introduced by directly applying MSQ. This transformation reduces memory consumption and improves efficiency of GEMM kernel, which would otherwise be compromised by the interleaved and non-fixed positions of visual and textual tokens. As shown in Table 17, MQuant can achieve up to 24.7% speedup and 152.9% memory savings.
A.11. Comparison of Different MLLMs: Input Pre-Process, LayerNorm Architecture in Vision Encoder, Model Parameters and Flops.
Table 19. Comparison of TTFT sensitivity to image resolution, model Parameters and flops in mainstream MLLMs. means the Flops values are measured with the number of visual tokens is 256.
MLLMs TTFT’s sensitivity to input image resolution LayerNorm Params (B) FLOPs (T)
Visual LLM Visual LLM
InternVL2-8B TTFT increases with input image aspect ratio Pre-LN 0.34 7.74 1.28 7.54
Qwen-VL-Chat-9.6B Fixed input resolution (448×448) Pre-LN 1.94 7.73 4.23 3.70
MiniCPM-V 2.6-8B TTFT increases with input image aspect ratio Pre-LN 0.49 7.61 4.16 3.64
Qwen2-VL-7B TTFT increases quadratically with input image resolution Pre-LN 0.68 7.61 1.31 3.61
GLM-4V-9B Fixed input resolution (1120×1120) Post-LN 4.51 8.78 12.10 4.70
In this section, we compare the visual input pre-process methods, LayerNorm structures in MLLM vision encoder, model parameters, and Flops across five mainstream MLLMs: InternVL2-8B (Chen et al., 2024b), Qwen-VL-Chat-9.6B (Bai et al., 2023b), MiniCPM-V 2.6-8B (Yao et al., 2024), Qwen2-VL-7B (Wang et al., 2024a), and GLM-4V-9B (Hong et al., 2024). As shown in Table 19, this comparison highlights the architectural differences between these models, particularly the TTFT sensitivity to input image resolution, illustrating the unique challenges these variations present for quantization and efficient inference.
A.12. Quantization Granularity
Furthermore, as mentioned in SoomthQuant  (Xiao et al., 2022), there are different different granularity levels. The per-tensor static quantization uses a single step size for the entire matrix. Per-token dynamic quantization employs different $s$ for the activations associated with each token, being a common granularity for activations quantization of existing LLMs. For weights, per-channel quantization applies distinct $s$ for each output channel of the weights, while group-wise quantization utilizes a coarse-grained $s$ for different channel groups. Notably, group-wise quantization is a prevalent granularity for weight quantization in LLMs  (Frantar et al., 2022).
A.13. LayerNorm, RMSNorm and Computational Invariance
We introduce LayerNorm, RMSNorm, computational Invariance, and their usage in Transformers.
Layer Normalization (LayerNorm, LN) (Ba, 2016) is a technique to normalize the activations of intermediate layers of neural networks. Given a vector $\mathbf{x}\in\mathbb{R}^{d}$, LayerNorm normalizes it to obtain a zero-mean unit-variance vector,
(17) $\text{LayerNorm}(\mathbf{x})=\frac{\mathbf{x}-\mu(\mathbf{x})\mathbf{1}}{\sqrt{\lVert{\mathbf{x}}\lVert_{2}^{2}/d-\mu^{2}(\mathbf{x})+\epsilon}},\text{where }\mu(\mathbf{x})=\frac{\mathbf{1}^{T}\mathbf{x}}{d},\epsilon>0.$
LayerNorm recenters and rescales the activations and gradients in the forward and backward computations, which enables fast and robust training of neural networks.
Root Mean Square Normalization (RMSNorm) (Zhang et al., 2019) is another technique used for normalizing the activations. It is similar to LayerNorm in that it aims to accelerate and stabilize the training but uses a different normalization approach. Instead of normalizing the inputs based on their mean and variance, RMSNorm normalizes them based on their root mean square (RMS) value. It is defined in the following equation,
(18) $\text{RMSNorm}(\mathbf{x})=\frac{\mathbf{x}}{\sqrt{\lVert{\mathbf{x}}\lVert_{2}^{2}/d+\epsilon}},\text{where }\epsilon>0.$
RMSNorm only rescales the input vector and the corresponding gradients, discarding the recentering process. As shown in their definitions, RMSNorm is computationally simpler and more efficient than LayerNorm. It is reported that replacing LayerNorm with RMSNorm can achieve comparable performance and save training and inference time by $7\%-64\%$ (Zhang et al., 2019).
Given a zero-mean vector $\mathbf{x}$, these two kinds of normalization are equivalent. Formally, if $\mu(\mathbf{x})=0$, then $\text{LayerNorm}(\mathbf{x})=\text{RMSNorm}(\mathbf{x})$. We may optionally introduce learnable parameters and apply an element-wise affine transformation on the output of LayerNorm and RMSNorm.
LayerNorm (LN) and RMSNorm
Given the input concated token $\bm{X}$ after embeddings with the shape $L\times D$, the $\bm{X}$ is passed through a LayerNorm (Ba, 2016) operation, which subtracts the mean from each row of the matrix, divides the row by its standard deviation, rescales (columnwise), and adds an offset. Follow  (Ashkboos et al., 2024a), we write the LayerNorm block as
(19) $\textrm{LayerNorm}(\bm{X})=\textrm{RMSNorm}(\bm{X}\bm{M})\textrm{diag}(\bm{\alpha})\sqrt{D}+\mathbf{1}_{N}\bm{\beta}^{\top}$
where $\textrm{RMSNorm}(\bm{X})$ applies $\mathbf{x}\leftarrow\mathbf{x}/\|\mathbf{x}\|$ to each row of $\bm{X}$, and $\bm{X}=concat(\bm{E}_{v},\bm{E}_{t}))$ is the concatenation between text tokens $\bm{E}_{t}$ and the visual tokens $\bm{E}_{v}$. The vector parameter $\bm{\alpha}$ and offset (vector) parameter $\bm{\beta}$ are learned independently at each LayerNorm instance. The constant matrix $\bm{M}=\bm{I}-\frac{1}{D}\mathbf{1}\mathbf{1}^{\top}$ is a $D\times D$ matrix which subtracts the mean from each row of $\bm{X}$, called recentering operation. Formally, if $\bm{M}=\bm{I}$, the input $\bm{X}$ has a zero-mean, the Eq 19 is equivalent to RMSNorm. Specifically, LayerNorm is widely employed in visual encoders $E$, whereas RMSNorm  (Zhang et al., 2019) is commonly used in LLMs (Touvron et al., 2023a; Dubey et al., 2024) and has been shown to accelerate training and inference time with similar performance (Zhang et al., 2019).
Computational Invariance in RMSNorm.
Based on the computational invariance, recent studies (Ashkboos et al., 2024a, b) have shown that orthogonal transformations can effectively smooth outliers and improve the quantize-ability of both weights and activations. In particular, for transformers, inserting linear layers with an orthogonal matrices $\bm{Q}$ before and after the RMSNorm (Zhang et al., 2019) layer in a transformer, the network remains unchanged. In detail, given the input $\bm{X}$ and orthogonal matrix $\bm{Q}$ for RMSNorm layer, the computational invariance means: $\textrm{RMSNorm}(\bm{X}\bm{Q})\bm{Q}^{\top}=\textrm{RMSNorm}$. Here, $\bm{Q}^{\top}\bm{Q}=\bm{Q}\bm{Q}^{\top}=\bm{I}$ and a rotation matrix is an orthogonal matrix with $|\bm{Q}|=1$. Note that multiplying a vector $\mathbf{x}$ by $\bm{Q}$ does not change the norm of the vector, since $\|\bm{Q}\mathbf{x}\|=\sqrt{\mathbf{x}^{\top}\bm{Q}^{\top}\bm{Q}\mathbf{x}}=\sqrt{\mathbf{x}^{\top}\mathbf{x}}=\|\mathbf{x}\|$.
A.14. LayerNorm to RMSNorm Transformation.
Figure 10. The proposed Post-LN + Rotate Scheme.
Post-LayerNorm to RMSNorm Transformation.
As shown in Figure 10, we present the detailed Post-LN + Rotate design. Unlike the Pre-LN + Rotate in SliceGPT (Ashkboos et al., 2024a), our Post-LN + Rotate scheme extends the applicability for different MLLMs’ vision encoder.
Pre-LayerNorm to RMSNorm Transformation.
Here, we propose unified LayerNorm-to-RMSNorm transformation, aiming to synthesize the transformer architecture of MLLMs’ vision encoders and LLMs, endowing them with rotation-friendly characteristics that facilitate the effective removal of outliers. We take Pre-LN transformer as an example to show that how transform Pre-LN into RMSNorm layer while guaranting arithmetic equivalence. As shown in Figure 11 (a), for the input $\bm{X}_{k}$ of the $k$-th block, the main block in pre-LN transformer is $\bm{X}_{k+1}=\bm{X}_{k}+\ell_{2}(g(\ell_{1}(LN(\bm{X}_{k}))))$, where $k\in[1,N]$, and $N$ is the block number. If $g$ is an activation function, such as GELU, this block is a multi-layer perceptron (MLP) module. If $g$ is a multi-head attention, then this block is the casual attention module (Vaswani et al., 2017). Due to the recentering operation, LN exhibits invariance to shifts, such that $LN(\bm{X}_{k}-a\mathbf{1})=LN(\bm{X}_{k}),\forall a\in\mathbb{R}$.
Figure 11. The illustration of transformation from Pre-LN to RMSNorm.
Therefore, as shown in Figure 11 (b), we can replace LN as RMSNorm layer through two modifications: recenter the input $\bm{X}_{k}$ to $\bm{X}_{k}-\mu(\bm{X}_{k})\mathbf{1}$, ensuring that the input to norm layer maintain a zero mean. adjust the weights $\bm{A}_{2}$ and bias $\mathbf{b}_{2}$ of the the linear $\ell_{2}$ to $\hat{\bm{A}_{2}}=\bm{A}_{2}-\frac{1}{D}\mathbf{1}\mathbf{1}^{T}\bm{A}_{2},\hat{\mathbf{b}_{2}}=\mathbf{b}_{2}-\mu(\mathbf{b}_{2})\mathbf{1}$. Consequently, the LN can be replaced with an RMSNorm layer with the same arithmetic functionality. The first operation is to recenter $\bm{X}_{k}$, while the second operation is to recenter the output of main branches. Notably, since $\bm{X}_{k+1}=\bm{X}_{k}+\ell_{2}(g(\ell_{1}(LN(\bm{X}_{k}))))$, after applying and , the input and the output of main branch are re-centered with zero-mean, while the input of residual branches also maintain a zero mean. Therefore, the output after current blocks, $\bm{X}_{k+1}$ (which serves as the input for next block), still maintain zero-mean. A detailed proof is provided in the Appendix. Ultimately, we establish the equivalence of Pre-LN and Pre-RMSNorm Transformers. Now that every LayerNorm in the transformer has been converted to RMSNorm in MLLMs, we can use any orthogonal matrices $\bm{Q}$ to the model. Therefore, the visual encoder and LLMs are in a rotation-friendly RMSNorm-only transformer architecture.
Figure 12. Illustration the weight distributions for the down-proj layer in different MLLMs’s visual encoder under three conditions: (a) original weight, (b) weights after FHT, and (c) weight after FHT with our RMS.
A.15. Weight Distribution Visualization with RMS.
To facilitate a better analysis of RMS, we also visualize weight distribution across other mainstream MLLMs (InternVL, Qwen-VL, MiniCPM-V and ChatVLM2). As shown in Figure  12, we can observe that the FHT-induced weight outliers presents at all MLLMs. The weight distribution reveals that FHT amplification increases weight magnitudes, while RMS effectively reduces these magnitudes, significantly improving quantization stability.
A.16. Discussion About Video-based MLLMs.
In video-based MLLMs like Video-LLaMA (Zhang et al., 2023) and Qwen2-VL (Wang et al., 2024a), the pipeline involves sampling video frames at fixed intervals, extracting visual features per frame using a vision encoder, and modeling temporal relationships with downstream components (e.g., via temporal positional encoding in Video-LLaMA or 3D convolution in Qwen2-VL). Two key features emerge: (1) videos are processed as multi-frame image sequences, and (2) Our Modality-specific static quantization is not affected by the temporal-aware module. Therefore, for an input video, MQuant’s per-modality static quantization applies naturally: we compute quantization parameters offline across these visual frames. Specifically, the dynamic nature of visual tokens is handled as an image sequence, not altering our static quantization’s efficiency or accuracy in MQuant (Section 3). We recognize this as a promising direction for future work.